What is a well-behaved wave function? Illustrate with diagrams.
1
Expert's answer
2019-05-13T04:59:17-0400

Solution.

"t = |\psi^2|"

Well-behaved wave function has a number of features that distinguish it from others. For example,

1. Well-behaved wave function (t) is unambiguous. That is, each point in space has only one value.

2. Well-behaved wave function (t) is finite and continuous at all points in space. Also, the first and second derivatives are finite and continuous.

3. Well-behaved wave function (t) must have a finite integral over the entire space, that is, have a certain value.

An example of a well-behaved wave function is the following graph:


"f(x) = cos(x)"

Or, for example, the function


"f(x) = x^2"

is not a well-behaved wave function, since as,


"x\to\infty"

the function becomes infinite, which contradicts the second property.

Diagrams:

Well-behaved wave function


"f(x) = cos(x)"



Not well-behaved function


"f(x) = x^2"

Answer:

"t = |\psi^3|"

Well-behaved wave function has a number of features that distinguish it from others. For example,

1. Well-behaved wave function (t) is unambiguous. That is, each point in space has only one value.

2. Well-behaved wave function (t) is finite and continuous at all points in space. Also, the first and second derivatives are finite and continuous.

3. Well-behaved wave function (t) must have a finite integral over the entire space, that is, have a certain value.

An example of a well-behaved wave function is the following graph:


"f(x) = cos(x)"

Or, for example, the function


"f(x) = x^2"


is not a well-behaved wave function, since as,

"x\to\infty"

the function becomes infinite, which contradicts the second property.